The length of a rectangle is 5cm longer than twice the length of the width. If the perimeter of the rectangle is 88cm, what is the width?
Answer by Boreal(15235)

width=x
length=2x+5
perimeter=88=2L+2W
so half perimeter=44=L+W
x+2x+5=44
3x=39
x=13 cm width ANSWER
2x+5=31 cm length